### v2

# Picogram-Level Nanoplastic Analysis with Nanoelectromechanical System Fourier Transform Infrared Spectroscopy: NEMS-FTIR

Abstract
We present a photothermal infrared spectroscopy-based approach for the chemical characterization and quantification of nanoplastics. By combining the high sensitivity of nanoelectromechanical systems (NEMS) with the wide spectral range and ubiquity of commercially available Fourier transform infrared (FTIR) spectrometers, NEMS–FTIR offers a cost-efficient and cryogen-free option for the rapid, routine analysis of nanoplastics in aqueous samples. Polypropylene, polystyrene, and polyvinyl chloride nanoplastics with nominal diameters ranging from 54 to 262 nm were analyzed by NEMS–FTIR with limits of detection ranging from 102 pg to 355 pg. The absorptance measured by NEMS–FTIR could be further converted to absolute sample mass using the attenuation coefficient, as demonstrated for polystyrene. Thanks to the wide spectral range of NEMS-–FTIR, nanoplastic particles from different polymers could be readily identified, even when present in a mixture. The potential of NEMS–FTIR for the analysis of real samples was demonstrated by identifying the presence of nanoplastics released in water during tea brewing. Polyamide leachates were identified in the brewing water without sample pre–concentration. Accelerated aging of the tea bags under elevated temperature and UV radiation showed further release of polyamide over time.
Keywords: nanoplastics, nanoparticles, NEMS–FTIR, Fourier transform infrared spectroscopy, photothermal sensing, nanoelectromechanical system, nanomechanical resonator, tea bag.
INTRODUCTION
Nanoplastics have become ubiquitous and pose significant environmental and health risks due to their high reactivity, potential to transport pollutants, and widespread dispersion across ecosystems [1, 2, 3, 4, 5, 6, 7]. The ability of nanoplastics to penetrate deep into tissues underscores the need for their routine chemical characterization and monitoring [8]. Sampling of nanoplastic particles is particularly challenging due to their nanoscale dimensions and low concentrations, which often exceed the capabilities of many analytical techniques [9, 10, 11]. While state–of–the–art techniques allow for routine microplastic detection and identification, nanoplastics, on the other hand, remain a challenge to this day.
Mass spectrometry–based techniques can be used for the bulk chemical identification of nanoplastics [10, 11, 12]. Pyrolysis–gas chromatography–mass spectrometry (Py–GC/MS) has been extensively applied to detect various types of nanoplastics, including polystyrene (PS), polypropylene (PP), polycarbonate (PC), polyethylene terephthalate (PET), polymethyl methacrylate (PMMA), polyvinyl chloride (PVC), polyethylene (PE), and Nylon 6/66, with reported limits of detection (LoDs) as low as 1 ng/L [13, 14]. Using the same technique, Li et al[15] estimated an LoD of 10 ng for two polymer types of nanoplastics in water, PS, and PMMA. A combination of atomic force microscope–IR (AFM–IR) spectroscopy and Py–GC/MS was also used to detect and quantify PE and PVC particles (20-–1000 nm), with LoDs of 0.28 $\mu$g for PE and 1.54 $\mu$g for PVC [16]. Alternative approaches, such as thermal desorption–proton transfer reaction-mass spectrometry (TD–PTR–MS), demonstrated remarkable sensitivity, with Materić et al[17] detecting pure 1 $\mu$m PS particles with an LoD of 0.34 ng, allowing detection of these particles in environmental samples as small as 1 mL without preconcentration. While mass spectrometry–based techniques provide excellent capabilities, their complexity and cost hinder their use in routine environmental monitoring [11, 12, 18].
FTIR spectroscopy with its broad spectral range allows functional group chemical identification of unknown samples and is one of the most extensively used methods in microplastic analysis [19, 20]. Commonly used attenuated total reflectance–FTIR (ATR–FTIR) spectroscopy is a contact bulk analysis technique with a detection limit in the order of 10 ng [21]. FTIR microscopy methods, such as µ–FTIR scanning microscopy and focal plane array–FTIR (FPA–FTIR) microscopy, offer a contact–less FTIR solution, reaching LoDs as low as 450 pg when used with a cryogenically cooled MCT array [22]. However, diffraction limits the reliable analysis of single particles with sizes below 10 $\mu$m [19, 20].
Quantum cascade laser infrared (QCL–IR) microscopy provides higher sensitivity, albeit with a narrower spectral range and additional issues with coherence artifacts. Although QCLs cover the fingerprint region relevant for polymer identification and provide stronger spectral signals due to their high source intensity, they typically have higher noise levels and limited spectral coverage, missing informative regions such as the C–-H stretching vibrations around 3000 cm-1 and the low-wavenumber region around 700 cm-1. Full–range spectra, as enabled by FTIR, also allow for chemometric analysis and spectral deconvolution, which can help identify and quantify nanoplastics in multicomponent samples. While such methods can also work within narrower spectral windows, having access to broader spectral information can make differentiation easier and more reliable. QCL–IR microscopy is a complementary method to FTIR microscopy and is therefore sometimes offered in tandem. [^2] HYPERION II from Bruker Corporation All standard IR spectroscopy techniques suffer from Mie scattering–related spectral artifacts when measuring particulate samples.[23]
Scattering, diffraction, and coherence issues can be mitigated by QCL–based photothermal scanning methods, such as optical–-photothermal IR (O-PTIR) spectroscopy or AFM-IR microscopy. O–PTIR has proven effective for the imaging with sub–micrometer resolution of individual PS nanoparticles as small as 250 nm in mammalian tissues [24] and identification of plastics (> 600 nm) containing polysiloxanes and imides, originating from the steam disinfection of silicone baby teats [25]. Even better lateral resolution and single nanoparticle analysis can be obtained by AFM-IR [26]. However, AFM–IR faces challenges with low signal–to–noise ratio (SNR) and slow imaging speed [19].
Raman spectroscopy has the advantage of probing at visible wavelengths with low interference from water and, similar to O–PTIR, allows for the analysis of plastic particles down to sub-micrometer sizes, with spatial resolution sufficient to resolve and chemically identify individual particles. [27]. Surface–enhanced Raman spectroscopy (SERS) has emerged as a promising technique. While Zhou et al[28] and Hu et al[29] reported LoDs as low as 5 $\mu$g/mL for PS particles as small as 50 nm using conventional SERS substrates, integration with sample preparation methods has pushed the sensitivity down to $10^{-4}$ $\mu$g/mL [30]. Hyperspectral stimulated Raman scattering (SRS) microscopy has been demonstrated as a powerful tool for the chemical imaging of nanoplastics, enabling identification and classification of particles down to 130 nm [31]. However, SERS suffers from issues such as spectral artifacts and reliance on carefully engineered substrates, limiting its versatility [11, 32]. SRS microscopy, despite offering rapid detection and high sensitivity, relies on complex and expensive instrumentation [33].
[…]
Although existing analytical techniques such as Py–GC/MS, QCL–IR, O–PTIR, or SERS offer high sensitivity and specificity, they remain impractical for routine environmental monitoring, particularly for water suppliers and official control laboratories [34].
Nanoelectromechanical systems IR spectroscopy (NEMS–IR) is a technique based on the photothermal effect that has been introduced in recent years. It employs NEMS membrane resonators, or chips, as its key component [35, 36]. The NEMS chip can be used simultaneously as a sample carrier and detector. The NEMS chips, typically consisting of a pre–stressed few–nanometer–thin silicon nitride (SiN) membrane, are robust and allow for a wide variety of sampling techniques [37], including direct aerosol collection [38, 36, 35] and drop casting. When a sample deposited on the NEMS chip’s surface absorbs IR light, local heating occurs, causing thermal expansion and tensile stress reduction. The corresponding frequency detuning of the NEMS chip is proportional to the absorbed power, and the frequency shift can be monitored using a closed–loop oscillation scheme in combination with a frequency counter, as demonstrated by Bešić et al[39, 40]. NEMS–IR has shown picogram sensitivity at room temperature over a large spectral range, from the ultraviolet to the far–IR [37]. This technique has been used to analyze polymer nanoparticles [41, 42], pharmaceutical compounds [36, 43], polymer micelles [44], thin polymer films [21] and explosives [45]. NEMS–IR with in–situ thermal desorption (TD) analysis has also been demonstrated for separating simple analyte mixtures by Luhmann et al[35].
NEMS–IR has only been used for dispersive spectroscopy, typically relying on narrow-band tunable QCLs. This paper introduces NEMS–FTIR spectroscopy (see Figure 1) as a new approach to the chemical characterization and quantification of nanoparticles, demonstrated here with nanoplastics. By combining the broad spectral coverage, wide availability, and ease of use of commercially available FTIR spectrometers with the high sensitivity of nanomechanical photothermal detection using NEMS membrane resonators, NEMS-–FTIR enables artifact–free measurements with picogram–level detection limits — all without the need for cryogenic cooling. Thanks to its straightforward sampling of nanoparticles dispersions directly onto the disposable NEMS chip, NEMS-–FTIR provides a powerful bulk analysis method for the routine detection and characterization of nanoplastics. Optomechanical metalized cantilevers have previously been used as photothermal detectors in combination with FTIR sources [46], but their lower sensitivity [47], reliance on optical readout, and limited surface area for sample collection have constrained their practical application.
Here we demonstrate the capability of NEMS–FTIR for the sampling, identification, and quantification of nanoplastics dispersions and nylon-based polyamide (PA) tea bag leachates. Many studies have analyzed the release of micro- and nanoplastic particles from tea bags during tea preparation using different analytical methods such as scanning transmission X–ray microscopy [48], IR photothermal heterodyne microscopy [49], Raman imaging [50], and ATR–FTIR spectroscopy [51, 52]. Most of these, however, require a much higher tea bag–to–water ratio than that applied in this study [52, 51] or involve sample up–concentration steps such as ultracentrifugation [51], density separation [48], or evaporation drying to obtain a powder [52]. In contrast, we demonstrate here that NEMS-–FTIR can chemically identify PA leachates in the brewing water from a single tea bag. NEMS-FTIR analysis of the leachate required just a few nL sample aliquot and did not require any pre–concentration steps.
RESULTS AND DISCUSSION
Single nanoplastic dispersions
Aqueous dispersions of three different polymers, PS (nominal diameter: 100 nm), PP (nominal diameter: 54 nm), and PVC (nominal diameter: 262 nm), were deposited via drop casting on individual NEMS chips. Figure 2A–C shows PS, PP, and PVC nanoparticles after deposition onto the central circular perforation of the NEMS chips (see Figure 1A). These scanning electron microscopy (SEM) images highlight the generally spherical morphology and relatively uniform size distributions of the nanoplastic particles studied. However, there are areas of nanoparticle agglomerations on the membrane of the NEMS chip caused by the coffee ring effect during droplet drying (see Figure S1 in Supporting Information). After drying, nanoparticles were confined within the perforated area of all analyzed NEMS chips, except in a few instances where minor droplet spreading occurred beyond this region. The spread–out particles still remained located within the confines of the NEMS membrane, which is the active area relevant for signal detection.
Figure 2: Characterization and quantification of nanoplastics. (A–C) SEM images of PS, PP, and PVC nanoparticles, respectively. (D) NEMS–FTIR spectra of varying mass loads of PS nanoparticles deposited on the NEMS chips. The inset highlights the 1452 cm-1 peak, which was used to construct the calibration curve and determine the LoD. (E) Calibration curves for the 1452 $\text{cm}^{-1}$ PS peak, 1377 $\text{cm}^{-1}$ PP peak, and 1427 $\text{cm}^{-1}$ PVC peak, with the inset showing a zoomed–in view of the region corresponding to lower mass loads (error bars N=3, 95% C.L.).
Spectral data were obtained by depositing a 20 nL drop of various nanoplastic concentrations on the NEMS chips (N=3 for each concentration). The NEMS–FTIR spectra of PS nanoplastic particles for each concentration, power–corrected, normalized using the SiN peak at 835 cm-1, and converted into absorbance, are shown in Figure 2D. The broad SiN peak, originating from the membrane material and spanning the 1000–700 cm-1 region, serves as a useful internal standard for spectral normalization. The details of the spectral processing procedure are described in "Processing of NEMS–FTIR spectra" in the Materials and Methods section. The spectra of PP and PVC nanoparticles are available in the Supporting Information in Figures S2 and S3, respectively.
The IR peaks in the NEMS–FTIR spectra of PS nanoparticles are in agreement with data reported in the literature [53, 54, 55, 56]. Notably, PS is characterized by peaks at 753 and 700 $\text{cm}^{-1}$, as well as by a mono–substituted benzene ring, as evidenced by a series of weak overtone and combination bands between 2000 and 1650 $\text{cm}^{-1}$ (so–called ’benzene fingers’), often difficult to detect in the spectrum if the concentration of molecules containing the benzene ring is low [55]. However, these peaks are clearly visible in the NEMS–FTIR spectra of PS, even for the very low masses in the nanogram range.
The NEMS–FTIR spectra of PP and PVC nanoparticles also display peaks which are in agreement with prior studies using FTIR spectroscopy [55, 57], as well as with the measurements obtained using early IR spectrophotometers employing prisms [58]. NEMS–FTIR does not suffer from the red shift typically associated with measurements performed by ATR–FTIR, as observed in previous studies [59]. In ATR–FTIR, slight shifts in some peak positions are observed due to the interaction of the evanescent wave with the sample surface, which depends on both the refractive index and wavelength. The wavelength-dependent penetration depth inherent to ATR measurements can also cause lower–wavenumber bands to appear more intense compared to transmission (or absorbance) spectra. [60, 61, 62, 63, 64]. This artefact is also absent from NEMS–FTIR spectra. ATR–FTIR performance is inherently constrained by the choice of crystal material, as different sample types require different ATR crystals for optimal spectral response [63]. This issue also does not apply to NEMS-FTIR, which relies on direct absorption. NEMS–FTIR generates spectra comparable to those obtained with transmission–FTIR spectroscopy, allowing seamless integration with existing spectral libraries and chemometric tools for efficient and accurate analysis. Furthermore, since the nanoplastic particles are analyzed under high vacuum, the resulting spectra are free from spectral interferences from water molecules and carbon dioxide (CO2).
Figure 2E shows the calibration curves for PS, PP, and PVC nanoparticles. Calibration curves were constructed using the characteristic peaks at 1452 $\text{cm}^{-1}$ for PS (see inset in Figure 2D), 1377 $\text{cm}^{-1}$ for PP (see inset in Figure S2), and 1427 $\text{cm}^{-1}$ for PVC (see inset in Figure S3) and the results showed good linearity ($\text{R}^{2}>0.9885$). LoDs were calculated from the calibration curves using the relation $3\sigma/m$, where $\sigma$ is the standard deviation of method blanks ($\text{N}=9$) and $m$ is the slope of the corresponding calibration curve. The results are summarized in Table 1. PP exhibited the steepest calibration slope, resulting from the strongest IR band used for the calibration curve, and consequently had the lowest LoD. Detection limits are lower than those typically achieved by Py–GC/MS [15] and are comparable to the performance of state–of–the–art TD–PTR–MS techniques [17]. The variations in detection limits for different nanoplastic particles did not correlate with nanoparticle size. In the measurements conducted on the chips with higher loadings, the observed nanoparticle agglomerations did not result in deviations in linearity.
The mass of PS nanoparticles present on the NEMS chips was estimated using the attenuation coefficient of PS and the absorptances calculated from the NEMS–FTIR spectra at $\tilde{\nu}=1452\text{ cm}^{-1}$. The mass estimation procedure is detailed in the "Mass estimation" in the Methods section. The calculation of the sample absorptance accounted for the spatial distribution of the particles, which predominantly formed a circular patch in the center of the perforation on the NEMS chips. Figure 3A shows the correlation between the deposited (amount of deposited PS particles) and estimated (from the measured absorptances of PS particles) mass of nanoplastic particles. The uncertainty on the sample mass deposited on the NEMS chip was estimated from the uncertainty on the tools used for sample preparation and for sample deposition. The largest uncertainty on the sample mass as estimated from the NEMS–FTIR spectra arises from the variability in the measured absorptance $\alpha_{\text{S}}(1452\leavevmode\nobreak\ \text{cm}^{-1})$. Exact values for the masses deposited on the NEMS chips and estimated masses as calculated from the NEMS–FTIR spectra, along with their standard deviations, can be found in Table S1 in the Supporting Information.
Figure 3B presents a comparison between the NEMS-FTIR spectrum of 20 ng of PS, expressed in absorbance, and a reference spectrum calculated from refractive index data [65] for an equal sample mass and spot size, showing agreement for both peak positions and intensity.
Table 1: LoD calculation parameters. Summary of the calculated LoDs for different nanoplastics, including the corresponding peak position (in wavenumbers), standard deviations of the blank ($\sigma$) at those respective peak positions (N = 9), and calibration slopes ($m$). LoDs were determined for $3\sigma/m$.
Plastic Type Peak position (cm-1) Standard deviation of the blank ( $10^{-4}$ AU) Slope ($10^{-4}$ AU/ng) LoD (pg)
PS 1452 0.50 4.25 353
PP 1377 0.30 8.84 102
PVC 1427 0.49 4.17 355
Figure 3: Quantitative analysis. (A) Correlation between the mass of PS nanoplastic particles deposited on the NEMS chip and the estimated mass derived from the measured NEMS–FTIR spectra after conversion to optical absorbance (N=3). The inset highlights the lower mass range for improved visibility. Error bars represent propagated uncertainties from both the measurement and estimation processes. The uncertainty in the IR beam spot size was not included in this analysis and could result in additional error. (B) Absorbance spectrum of 20 ng PS measured via NEMS–FTIR compared to a reference spectrum calculated with Eq. (4) and (7) from PS refractive index data [65].
While NEMS–FTIR can be used to detect very small masses of analytes, very dilute solutions do require pre–concentration. The largest droplet volume that can be drop cast on a NEMS chip is limited to approximately 500 nL. Considering an LOD of 353 pg for PS (see Table 1), the minimum detectable concentration of PS is 0.7 $\mu$g/mL for a single 500 nL drop of sample. However, in addition to usual pre–concentration techniques such as solid phase extraction, ultrafiltration, and diafiltration, the NEMS–FTIR chip allows new sample pre–concentration possibilities. Multiple droplets can be deposited consecutively on the NEMS chip with intermediate drying steps to increase the analyte mass deposited to the desired level. Sampling via nebulization can also be used to concentrate larger sample volumes onto the NEMS chip. Based on a typical nebulizer flow rate of 20 $\mu$L/min, a sampling time of 1 hour, and an aerosol capture efficiency of approximately 50 % for 100 nm particles [38], we estimate that concentrations as low as 0.6 ng/mL could be measurable. When dealing with complex matrices containing high concentrations of organic material relative to the target analyte, matrix simplification or removal may be required to remove spectral interferences or avoid overloading the NEMS chip.
Nanoplastic mixtures
The SEM images in Figure 4A, B provide the close–up views of the perforated area of the NEMS chip sampled with PS, PP, and PVC nanoparticles previously mixed in a 1:1:1 mass ratio.
The NEMS–FTIR spectrum of the mixture in Figure 4C shows the characteristic peaks of each polymer component, highlighted by vertical lines. Despite a deposited mass of only 5 ng per component, the peaks corresponding to PS, PP, and PVC are clearly distinguishable. The exact peak positions of the PS, PP, and PVC nanoparticles can be found in the Supporting Information, section A. In environmental contexts, where mixtures of different polymer micro- and nanoparticles are commonly encountered, the broad spectral range of NEMS–FTIR increases the likelihood of detecting distinct peaks without interference, facilitating the identification of individual components when their spectral features are sufficiently resolved.
Figure 4: Characterization of nanoplastic mixture. (A) SEM image of a mixture containing PS, PP, and PVC. (B) Magnified view of the membrane area highlighting PS (⌀ 100 nm, blue square), PP (⌀ 54 nm, green square), and PVC nanoparticles (⌀ 262 nm, red square). (C) Stacked NEMS–FTIR spectra of pure PS, PP, and PVC nanoparticles and their 1:1:1 mixture. Vertical lines mark the characteristic peaks of each polymer.
Tea bag leachates
Figure 5A shows the schematic of the experimental setup for the quantification of nylon tea bag leachates. The empty tea bag was immersed in 200 mL of water preheated at 95°$\degree$°C for 10 minutes.
Figure 5: Characterization of PA particles released from nylon tea bags. (A) Schematic illustration of the sample preparation process. (B) SEM image showing fragments released into water by a nylon tea bag. (C) NEMS–FTIR spectra of nylon tea bag leachates (TBL; 100 nL, and 500 nL aliquots deposited onto NEMS chips without pre–concentration) compared to ATR–FTIR spectrum of 500 nL of tea bag leachate (TBL) and the ATR–FTIR spectrum of the nylon tea bag (TB). Vertical lines indicate characteristic IR peaks associated with PA. (D) NEMS–FTIR spectra of the leachates from nylon tea bags subjected to accelerated weathering for 25 consecutive days, simulating one year of aging under adjusted conditions. The right inset graph highlights the characteristic PA peak at 1640 cm-1, corresponding to the amide I band. The left inset graph represents the dependence of signal intensity at 1640 cm-1 on the number of days during which the nylon tea bag was subjected to accelerated aging.
Figure 5B shows micro- and nanoplastic fragments released from the nylon tea bag into the water during standard tea preparation, which were deposited onto the NEMS chip via nanodrop casting. Smaller fragments were also observed on the membranes after sampling the water from the aged tea bags, where a different sample deposition method, based on aerosolization of the liquid, was used (see Figure S4).
Figure 5C shows the NEMS–FTIR spectra of nylon tea bag leachates prepared in triplicates at two aliquot volumes, 100 nL and 500 nL, deposited onto NEMS chips as-is, without prior concentration steps. A comparison with the ATR–FTIR spectrum of 500 nL of the same tea bag leachate reveals that this volume was insufficient for ATR–FTIR detection, as no characteristic peaks were identified. In contrast, the reference ATR–FTIR spectrum of the nylon tea bag itself exhibits well-defined spectral features that correspond closely to those observed in the NEMS–FTIR spectra of the leachates. The high sensitivity of NEMS–FTIR not only enables the use of very small sample volumes but also significantly reduces preparation time, as no concentration is necessary. Volumes of 500 nL of leachate were deposited on the NEMS chip and dried within approximately 30 minutes using the drop casting method (see Sampling methods in Methods). Since this volume failed to produce a detectable signal with ATR–FTIR, a larger sample would be required, which in turn implies longer drying times and more time-consuming sample preparation.
The raw, uncorrected ATR-FTIR spectrum of the tea bag is shown in Figure S5A, alongside the corrected spectrum to illustrate the effects of the applied automatic ATR correction procedures. As can be observed in Figure 5C (see closeup spectrum in Figure S5B), weak diamond phonon absorption bands originating from the ATR crystal [63] appear in the 2500–1800 cm-1 region. These arise due to the refractive index contrast between the sample and diamond ($n=2.42$ [66]). Since the refractive index of nylon-based PA is around 1.50–1.53 [67], the difference ($\Delta n$) is sufficient to make the crystal phonon features visible in the spectrum [63]. Additionally, CO2 absorption peaks are visible within the same region (Figure S5B), contributing to the observed spectral features. The variations in peak position and peak shapes, as well as the presence of additional diamond absorption bands associated with ATR-FTIR spectra, CO2 peaks and potential water vapour peaks, are all absent from the NEMS–FTIR spectra.
Vertical lines, as shown in Figure 5C, highlight characteristic IR peaks corresponding to the chemical structure of PA [68, 69]. The peak at 1720 cm-1 in the NEMS–FTIR spectrum likely originates from imide groups formed during thermal–oxidative degradation of nylon, induced by brewing the tea bag in 95°$\degree$°C water [70]. Exact peak positions can be found in the Supporting Information, section B. The same peaks appear in the NEMS–FTIR spectra of water samples in which individual empty nylon tea bags were immersed and subjected to accelerated aging conditions. As aging progressed, the nylon tea bags degraded further, releasing an increasing amount of PA particles, reflected in the peaks’ increasing intensity as shown in the insets (Figure 5D). This observation is in line with previous studies suggesting that prolonged exposure to environmental aging conditions contributes to the release of polymeric particles into the surrounding medium [71].
The NEMS–FTIR spectra collected in both cases allowed for chemical identification of the material from which the tea bags were made and confirmation that this material was leaching into the tea water (Figure 5C, D). Characteristic IR peaks were observed, matching well with the reference ATR–FTIR spectrum of the original tea bag, as well as with the literature [68, 69], confirming the presence of PA, including the amide I (1641 cm-1) and amide II (1553 cm-1) bands. The amide I band primarily arises from C=O stretching vibrations in the amide group, while the amide II band results from a combination of N–H bending and C–N stretching vibrations, which are distinctive features of PA materials. The comparison with online spectral databases, such as Open Specy [72] and Spectragryph [73], identified the released material as nylon–based PA.
CONCLUSIONS
NEMS-FTIR has proven effective for rapidly identifying and quantifying nanoplastics in aqueous dispersions at picogram-level sensitivity. It was demonstrated that drop-cast or nebulized polymer nanoparticles could be identified from polydisperse mixtures and that the measured signals from deposited PS nanoparticles could be used to estimate the mass of PS present on the NEMS chip.
To demonstrate applicability to real-world samples, NEMS-FTIR was used to identify polymer particles released from nylon tea bags during brewing. Leachates were directly analyzed on NEMS chips without prior sample preparation or preconcentration. Compared to ATR-FTIR, NEMS-FTIR offered significantly higher sensitivity, revealing distinct nylon spectral features from as little as 100 nL of sample, whereas ATR-FTIR failed to show peaks even at five times that sample volume. Accelerated aging experiments on nylon tea bags demonstrated the monitoring of the release of polymer particles under prolonged environmental stress. Unlike ATR-FTIR, which requires complex spectral data corrections, careful crystal selection, and suffers from spectral artifacts, NEMS-FTIR simplifies data interpretation through direct absorption measurements. Measurements performed under high vacuum with a nitrogen-purged optical path minimize spectral interference from ambient CO2 and water vapor, and the SiN NEMS chip material can be used as an intrinsic internal standard. NEMS-FTIR spectra are compatible with standard transmission FTIR spectral databases, facilitating the reliable identification of unknown samples.
Taking advantage of the wide availability of FTIR instrumentation, NEMS-FTIR is ideally suited for routine monitoring and widespread adoption. Its non-destructive nature further allows seamless subsequent analysis by complementary techniques such as SEM, EDX, or O-PTIR for detailed sample characterization. The demonstrated performance and ease of integration with existing workflows highlight NEMS-FTIR as a promising tool for environmental monitoring and nanomaterials analysis in general.
METHODS
Nanoplastic particles
PS nanoparticles (⌀100 $\pm$ 10 nm, Prod. No. 43302), as a 10 % w/v dispersion in water, were obtained from Sigma–Aldrich (MO, USA). PP (⌀54 nm, Prod. No. PP50) and PVC nanoparticles (⌀262 $\pm$ 30 nm, Prod. No. PVC250), both as 1 % w/v dispersion in water, were obtained from Lab261 (CA, USA).
Preparation of single and mixed nanoplastic dispersions
[…]
Serial dilutions and mixture preparation were carried out using UHPLC–MS grade water. A PIPETMAN P1000 micropipette, with matching DIAMOND D1000 TIPACK tips (Prod. No. F144059M and F171500, respectively, Gilson Incorporated, WI, USA), was used for the preparation. All dilutions were stored in hydrolytic class 3 soda-lime glass vials with snap–on lids (Prod. No. LC84.1 and LC87.1, respectively, Carl Roth GmbH, Germany).
All stock and diluted dispersions were first vortexed for 30 seconds using a Mini Vortexer (Heathrow Scientific, IL, USA) to ensure consistent distribution before proceeding with the successive dilution. PS nanoparticle dispersions were prepared at concentrations of 31.25-2000 $\mu$g/mL, while PP and PVC dispersions ranged from 31.25 to 500 $\mu$g/mL. The mixed dispersion of all three types of polymer nanoparticles was prepared in a 1:1:1 mass ratio (PS:PP:PVC). Table 2 summarizes all final concentrations used for each single and mixed sample. Blank samples, consisting of the same UHPLC–MS grade water used for dilution procedures, underwent the same preparation process as the experimental samples.
Table 2: Characterization of nanoplastic dispersions. Concentrations of dispersions (A) and total deposited masses of nanoplastic particles (B), differentiating between single- and three-component dispersions. Empty cells indicate concentrations not used in the experiment for the corresponding dispersion.
Nanoplastics type Concentration ($\mu$g/mL) (A)
2000 1000 500 250 125 62.5 31.25
Total deposited mass (ng) (B)
PS 40 20 10 5 2.5 1.25 0.625
PP 10 5 2.5 1.25 0.625
PVC 10 5 2.5 1.25 0.625
PS:PP:PVC (1:1:1) 5
Tea bags
Standard tea brewing procedure
The tea bags used in the experiment were purchased online (Temu.com, China), and the procedure for releasing plastic particles from the tea bags was adapted from the study by Hernandez et al. [52]. To minimize potential contamination, the tea bags, made of nylon according to the manufacturer, were purchased empty, eliminating the need for additional cutting or removing the tea leaves. The thin cotton string that attaches a tiny label to the tea bag was carefully removed with scissors. The tea bags were then rinsed five times with UHPLC–MS grade water to avoid impurities or loose plastic particles and allowed to dry overnight in a laminar flow hood.
To simulate a typical tea brewing process, 200 mL of UHPLC–MS grade water was poured into an Erlenmeyer flask (Duran, DWK Life Sciences GmbH, Germany), which had been pre–cleaned following the same procedure described in the "Preparation of single and mixed nanoplastic dispersions" section. Once the water reached its boiling point, the flask was removed from the hot plate and a single tea bag was immersed. After ten minutes, the tea bag was carefully removed using clean metal tweezers, and the water was allowed to cool overnight. The flask was covered with aluminum foil to prevent contamination during the cooling period. A blank sample was prepared using the same procedure with the same UHPLC–MS grade water but without the tea bag.
Accelerated aging of tea bags
The accelerated aging experiment was carried out in a SimTech Feutron Double Climate Chamber (73 x 77 x 102 cm, Feutron Klimasimulation GmbH, Germany) equipped with two Ultra–Vitalux lamps (300 W, Prod. No. 4008321543929, Osram, Germany) placed 30 cm apart (see Figure S6 in the Supporting Information). These lamps emitted broad-spectrum UV light (UVA: 315-400 nm, UVB: 280-315 nm) with an irradiance of approximately 59.1 W/m2, simulating natural solar radiation to induce degradation.
A metal plate, 32 cm below the lamps, held hydrolytic class 1 soda–lime glass vials with screw caps (Prod. No. LC92.1 and LE03.1, respectively, Carl Roth GmbH, Germany) containing the tea bag samples. The vials allowed good UV transmission, particularly in the UVA range, while minimizing light absorption.
The chamber was operated according to a cyclic schedule: 8 hours of UV exposure at 60°C, followed by 4 hours in the dark at 25°C, mimicking natural day and night temperature fluctuations. This regimen, based on ISO 4892-3 [74] and ASTM G154 [75] standards, lasted 25 days, which is equivalent to approximately one year of natural aging [76]. Seven PA tea bags, each in one glass vial with 30 mL of UHPLC–MS water, were artificially aged, and individual water samples were collected after 1, 2, 5, 10, 15, 20, and 25 day(s) to monitor particle release and degradation. Simultaneously, blanks (water without tea bags) were collected with each respective tea bag sample.
NEMS chips
As shown in Figure 1A, NEMS chips, or resonators, play a dual role as both sensing devices and sample holders, making them the core of NEMS–FTIR technology. The wafers used in this study to fabricate NEMS chips are made from silicon (Si) and SiN, with thicknesses of 380 $\mu$m and 50 nm, respectively, and an instrinsic tensile stress of approximately 50 MPa for the SiN. After the final manufacturing step, the KOH etching, a 1x1 mm2 SiN membrane is released. This membrane has a perforated area with a diameter of approximately 600 $\mu$m, consisting of 6 $\mu$m holes spaced 3 $\mu$m apart. It acts as the platform on which the sample is deposited for analysis. Two gold electrodes, each 10 $\mu$m wide, are positioned across the chip to enable electrical transduction, allowing mechanical vibrations to be converted into electrical signals, as demonstrated by Bešić et al[42] and Luhmann et al[35].
Sampling methods
Nanoliter droplet dispensing
Various drop casting strategies may be applied depending on the specific requirements of sample deposition. In this work, two different drop casting approaches were employed to deposit samples onto NEMS chips, selected based on the target deposition volume.
A PIPEJET nanoDispenser (BioFluidix-Hamilton Freiburg, Germany), shown in Figure S7A, was used to dispense the dispersions on the membrane of the NEMS chips. Polyimide capillaries with an inner diameter of 200 $\mu$m (Prod. No. PJ-20010) were used to deliver individual droplets of 20 nL. A 1 cm length of Tygon tubing (Prod. No. VM-20053-1) was attached to the capillary to facilitate sample loading. Liquid samples were inserted into the tubing using a PIPETMAN P1000 micropipette with matching DIAMOND D1000 TIPACK tips (Prod. No. F144059M and F171500, respectively, Gilson Incorporated, WI, USA). The stroke velocity was set to 90 $\mu$m/ms during droplet dispensing.
For each concentration of the dilution series, droplets of 20 nL were deposited onto the center of the individual NEMS chips. The total deposited mass of each sample is listed in Table 2B. Each deposition was performed in triplicate, with three NEMS chips prepared for each specific polymer mass. Nine additional NEMS chips were sampled with blank samples to provide a reliable baseline for subtraction.
For the mixture of nanoplastic particles, a single replicate was prepared by depositing one 20 nL drop of the pre–mixed dispersion onto a NEMS chip. Blank samples were deposited in parallel onto three additional NEMS chips.
For the analysis of the tea bag leachates, three NEMS chips were sampled with leachate samples and three with blank samples using the PIPEJET nanoDispenser. Each NEMS chip was sequentially loaded with ten drops (10 nL each) of liquid sample, with a two–minute delay between drops, totaling 100 nL per chip.
Furthermore, three NEMS chips were prepared using the Drop casting Accessory (Invisible–Light Labs GmbH, Austria), a PIPETMAN P2 micropipette, and corresponding DIAMOND DL10 TIPACK tips (Prod. No. F144054M and F171200, respectively, Gilson Incorporated, WI, USA). A single 500 nL droplet of the tea bag leachate was deposited onto each chip. The chips were then left to dry inside the Drop casting Accessory, shown in Figure S7B, in an environment with a controlled humidity gradient, which, due to the presence of perforations in the chip, facilitates droplet drying within the perforated area. In parallel, three additional NEMS chips were sampled using blank samples.
Aerosolization method for particle deposition
A system incorporating the Portable Aerosol Generation System (PAGS, Handix Scientific Inc., CO, USA) was employed to sample particles released from the nylon tea bags subjected to accelerated weathering cycles. This setup enabled consistent aerosolization and deposition of particles onto the NEMS chips facilitated by the holes in the central region of the membrane. Sampling was conducted at regular intervals (day 1, day 2, day 5, day 10, day 15, day 20, and day 25) to monitor the temporal evolution of particle release. The NEMS chip was placed in a holder (Aerosol Flow Adapter, Invisible–Light Labs GmbH, Austria), which was then positioned in an eight-channel filter sampler (FILT, Brechtel Manufacturing Inc., CA, USA) equipped with a built-in pump. The pump was set to a flow rate of 0.5 L/min for 5 minutes for each NEMS chip, which allowed sample deposition onto the membranes through the inertial impaction mechanism [38]. A schematic of the setup is presented in Figure S7C.
Experimental setup
SEM imaging
A Hitachi SU8030 SEM (Hitachi, Japan) was employed to confirm the sizes of the particles deposited on the NEMS chips and to analyze the leachates from the plastic tea bags. For imaging, the sampled NEMS chips were mounted on an aluminum sample holder. Secondary electron (SE) detection in the upper lens mode was utilized to capture high-resolution surface details. The microscope operated at an accelerating voltage of 2 kV with an emission current of 3900 nA. NEMS chips were examined at various magnifications to assess the particle dimensions and the structural characteristics of the deposited material.
NEMS–FTIR spectroscopy
[…]
During the measurement process, sampled NEMS chips were placed in the vacuum chamber of the nanoelectromechanical IR analyzer at 10-5 mbar and the temperature of the chips was consistently regulated to 25°$\degree$°C using the integrated thermoelectric cooling (TEC) controller and Peltier element. To minimize atmospheric interferences, the optical path of the EMILIE and FTIR spectrometer were purged with dry nitrogen throughout the measurements. A resolution of 4 cm-1, with a stabilization delay of 30 ms, 200 co-additions, and an aperture of 6 mm, were used for the FTIR parameters. The spectral range extended from 4000 to 400 cm-1.
ATR–FTIR spectroscopy
ATR–FTIR spectroscopy was used alongside the NEMS–FTIR to analyze the tea bag leachates as a reference IR technique. The tea bag material itself was also examined using ATR-FTIR to confirm it was the source of the detected components in the leachates. A Platinum ATR accessory with a diamond crystal (A225/Q, Bruker Optics, MA, USA) was used for all ATR measurements, performed on the same Bruker Vertex 70 spectrometer utilized for the NEMS–FTIR experiments.
To match the NEMS–FTIR experimental conditions, an aliquot of 500 nL of tea bag leachate was applied to the isopropanol–cleaned ATR diamond crystal. The droplet was dispensed using the same micropipette and tip combination as that used for sampling the NEMS chips with the same tea bag leachates (as described in the Nanoliter droplet dispensing section). Once the droplet had evaporated, the ATR measurement was performed by lowering the pressure clamp onto the remaining residue. Additionally, 500 nL of the corresponding blank sample was measured in the same manner.
A section of the tea bag material used in the brewing experiments was also analyzed. The cleaned ATR crystal was covered with the tea bag fragment, and the pressure clamp was lowered to ensure a good contact between the sample and the crystal.
Each measurement consisted of a background spectrum recorded with the empty crystal under pressure, followed by the sample spectrum (blank, leachate, or tea bag). All spectra were acquired with a resolution of 4 cm-1, over the spectral range of 4000-400 cm-1, using 32 scans, a scanner velocity of 2.5 kHz, and an aperture of 6 mm. Atmospheric compensation was enabled during all measurements to minimize the influence of the water vapor and CO2 absorption.
The ATR–FTIR spectra were processed using the "Extended ATR correction" (standard mode) and "Baseline correction" features in the OPUS software to correct for the slight distortions in peak intensity, shape, and position inherent to ATR-FTIR measurements [60, 61, 62, 63, 64]. The spectra were also smoothed using Savitzky-–Golay filtering (window length 20, polyorder 2).
Processing of NEMS-FTIR spectra
[…]
Spectral calibration was carried out using the intrinsic SiN absorption peak at 835 cm-1 as a reference. This calibration was based on transmission measurements through 12 clean, perforated, $2\times 2$ mm2, 50 nm thick low–stress SiN NEMS membranes. Figure S8 in the Supporting Information shows the average transmittance and corresponding absorptance of the NEMS chip’s SiN material measured using the internal detector of the FTIR spectrometer (2 mm aperture, 32 scans, and a resolution of 2 cm-1). Despite the different sizes of the membranes ($1\times 1$ mm2 and $2\times 2$ mm2) and perforation areas (600 $\mu$m and 1200 $\mu$m, respectively), the ratio of perforated to non–perforated membrane material remains consistent across both membrane types, making the calibration based on the SiN absorption peak possible.
The absorptance for a specific wavenumber $\tilde{\nu}$, $\alpha(\tilde{\nu})$, was calculated from the transmittance $T(\tilde{\nu})$ as $\alpha(\tilde{\nu})=1-T(\tilde{\nu})$, assuming insignificant scattering, yielding $\alpha_{\text{SiN}}(835\text{ cm}^{-1})=0.21\pm 0.01$. The NEMS–FTIR response of the bare SiN chip, $R_{\text{SiN}}$, was then related to absorptance through a calibration factor $\beta$:
$\beta=\frac{\alpha_{\text{SiN}}(835\text{ cm}^{-1})}{R_{\text{SiN}}(835\text{ cm}^{-1})}.$ (1)
Depositing the nanoplastics on the membrane with the nanodropper created a circular sample spot (see Figure S1 in Supporting Information). The nanoplastics occupied the area $\Sigma_{\text{S}}$, which is smaller than the area of the IR beam illuminating the chip, $\Sigma_{\text{IR}}$. Furthermore, the increased photothermal response, if the sample is concentrated in the membrane center instead of being evenly distributed over the entire membrane, was corrected by $\gamma$ [47]. The sample absorptance $\alpha_{\text{S}}(\tilde{\nu})$ can now be calculated from the measured NEMS-FTIR signal of the sample, $R_{\text{S}}$, as:
$\alpha_{\text{S}}(\tilde{\nu})=\beta R_{\text{S}}(\tilde{\nu})\frac{\Sigma_{\text{{IR}}}}{\Sigma_{\text{S}}}\frac{1}{\gamma}.$ (2)
[…]
$\Sigma_{\text{S}}$ corresponds to the diameter of the central nanoplastics area on the membrane with a diameter of $d_{\text{S}}=504\pm 46$ $\mu$m, based on six independent measurements. The total sample area is then given by $\Sigma_{\text{S}}=\eta\pi(d_{\text{S}}/2)^{2}$, with the fill factor of the perforation $\eta=0.62$.
$\gamma$ was calculated by finite element method (FEM) simulations by Kanellopulos et al[78]. The results for a square low-stress SiN membrane with a side length of 1 mm are given in Figure S9 in the Supporting Information, yielding $\gamma=1.68\pm 0.05$ for the given sample spot size. The nanoplastics spots on the membrane formed a coffee ring, as seen in (see Figure S1 in the Supporting Information). Up to 90% of particles typically accumulate at the droplet perimeter in the coffee ring, as shown for nanometer-sized PS dispersions [79, 80]. FEM simulations have shown that the formation of a coffee ring has no significant effect on the NEMS–FTIR response.
Finally, the blank-corrected nanoplastic absorbance can be calculated from
$A(\tilde{\nu})=-\log_{10}(1-[\alpha_{\text{S}}(\tilde{\nu})-\alpha_{\text{B}}(\tilde{\nu})]),$ (3)
with the absorptance of the blank $\alpha_{\text{B}}(\tilde{\nu})$.
Mass estimation
This framework aims to estimate the actual mass of the nanoplastic on the NEMS chip responsible for the NEMS-FTIR signal. Following the approach of Dudani and Takahama [81], it was assumed that particle scattering is negligible compared to absorption; thus, correction factors related to scattering are not required. Particles are much smaller than the wavelength of IR radiation used in this study, placing them in the Rayleigh regime, where absorption dominates IR attenuation. In this case, the absorbance $A(\tilde{\nu})$ of the deposited particles can be expressed as a function of the linear attenuation coefficient $\mu_{\text{10}}(\tilde{\nu})$ :
$A(\tilde{\nu})=\mu_{\text{10}}(\tilde{\nu})h_{\text{eff}}=\mu_{\text{10}}(\tilde{\nu})\dfrac{m}{\rho\Sigma_{\text{S}}},$ (4)
where the effective film thickness $h_{\text{eff}}$ is expressed as a function of the analyte mass $m$, the nanoparticle material density $\rho$, and the sample area $\Sigma_{\text{S}}$.
In NEMS–FTIR, the measured absorptance values are small$(\alpha_{\text{S}}(\tilde{\nu})\ll 1)$ and the absorbance (3) can be approximated using a first-order Taylor series expansion:
$A(\tilde{\nu})\approx\frac{\alpha_{\text{S}}(\tilde{\nu})-\alpha_{\text{B}}(\tilde{\nu})}{\ln(10)}.$ (5)
Combining Eqs. (2), (4), (5), and solving for the analyte mass gives:
$m=\beta\frac{R_{\text{S}}(\tilde{\nu})-R_{\text{B}}(\tilde{\nu})}{\mu_{\text{10}}(\tilde{\nu})}\frac{\rho}{\ln(10)}\frac{\Sigma_{\text{IR}}}{\gamma},$ (6)
where $R_{\text{B}}(\tilde{\nu})$ is the NEMS–FTIR response of the averaged blank spectra, and $\tilde{\nu}$ denotes the wavenumber corresponding to a characteristic vibrational mode of the sample.
The attenuation coefficients of the substance are determined from its complex refractive index, defined as $\tilde{n}=n+ik$, where $n$ and $k$ are the real and imaginary components, respectively. For a sparsely distributed particle layer, the decadic linear attenuation coefficient is given by:
$\mu_{\text{10}}(\tilde{\nu})=\dfrac{6\pi\tilde{\nu}}{\ln{(10)}}\operatorname{Im}\left\{\dfrac{\tilde{n}^{2}(\tilde{\nu})-1}{\tilde{n}^{2}(\tilde{\nu})+2}\right\}.$ (7)
A detailed derivation of Eq. (7) can be found in the Supporting Information, section C.
Supporting Information
This file includes:
Lists of characteristic peak positions (in wavenumbers) related to nanoplastic particles and tea bag leachates, derivation of the decadic linear attenuation coefficient for sparse particle film, SEM images of the coffee ring effect, NEMS-FTIR spectra of polypropylene and polyvinyl chloride nanoparticles, SEM images of micro- and nanofragments from tea bag leachates, comparison of non-corrected and corrected ATR-FTIR spectrum of nylon tea bag material, comparison between NEMS-FTIR and ATR-FTIR spectra regarding the spectral artifacts, experimental setup for the accelerated aging of the tea bags, sampling setups, IR spectral features of silicon nitride membranes, dependence of relative photothermal responsivity on the IR beam radius, values for deposited and estimated mass of polystyrene nanoparticles along with the standard deviations
Acknowledgements
The authors sincerely thank Jürgen Smoliner, Alois Lugstein and Masiar Sistani from the Institute of Solid State Electronics at TU Wien for letting us use their FTIR spectrometer. The authors gratefully acknowledge Noelia Barrabés Rabanal and the Research Group for Model Catalysis and Applied Catalysis at TU Wien for kindly providing access to their ATR module and supporting our measurements. The authors thank Michael Schneider and Paolo Martini from the Institute of Sensor and Actuator Systems at TU Wien for acquiring the SEM images and helping with finite element method simulations, respectively. Special thanks go to the Invisible-Light Labs team for their technical expertise and supportive encouragement, as well as to colleagues from the MNS group at TU Wien (Institute of Sensor and Actuator Systems) for their valuable insights and collaboration. The authors also gratefully acknowledge Georg Pfusterschmied for providing the image in Figure 1A. The authors gratefully thank Mihnea Surdu from the Extreme Environments Research Laboratory at EPFL in Switzerland for his assistance with the aerosol sampling setup.
Funding: This project received funding from the European Commission under grant agreement HORIZON–EIC–2021–TRANSITIONOPEN–101058011 – NEMILIES and 101202787 EMILIE Booster, the Austria Wirtschaftsservice (AWS SEED, Deep Tech & Physical Sciences project P2422873), the Austrian Science Fund (FWF) under project I 6086–N, and the Novo Nordisk Foundation under grant number NNF22OC0077964 – MASMONADE.
Author contributions:
Conceptualization: JTP, JPL, and SS. Methodology: JTP, JH, EŠ, NL, HB, JPL, and SS. Investigation: JTP, JH, EŠ, NL, AG, and HB. Supervision: AG, JPL, and SS. Writing-original draft: JTP and EŠ. Writing-review & editing: JTP, JPL, NL, and SS.
Competing interests: NL, HB, JPL, and SS are co–founders of Invisible–Light Labs GmbH who provided the analytical instrumentation used in this study. JTP, NL, and JH are employed by Invisible–Light Labs GmbH. EŠ and AG declare no competing interests.
Supporting Information for
Picogram-Level Nanoplastic Analysis with Nanoelectromechanical System Fourier Transform Infrared Spectroscopy: NEMS-FTIR
Jelena Timarac–Popović1,2*, Johannes Hiesberger1,2, Eldira Šesto1, Niklas Luhmann2,
Ariane Giesriegl1, Hajrudin Bešić1, Josiane P. Lafleur2, and Silvan Schmid1Correspondence e–mail address: jelena.popovic@tuwien.ac.at, silvan.schmid@tuwien.ac.at
1TU Wien, Institute of Sensor and Actuator Systems, Gusshausstrasse 27–29, 1040 Vienna, Austria.
2Invisible–Light Labs GmbH, Taubstummengasse 11, 1040 Vienna, Austria.
(Dated: July 19, 2025)
A.   NEMS–FTIR spectra of single NPs
The vertical lines observed in the NEMS–FTIR spectra (visible on Figure 4C) correspond to characteristic peak positions (in wavenumbers) associated with specific nanoplastic particles. Some lines correspond to potential minor oxidation or contamination during sample preparation.
The NEMS–FTIR spectrum of PS nanoparticles exhibit vertical lines at the following wavenumbers: 700, 753, 1452, 1493, 1601, 1745, 1800, 1868, 1943, 2849, 2923, 3026, 3060, and 3081 $\text{cm}^{-1}$.
The NEMS–FTIR spectrum of PP nanoparticles exhibit vertical lines at the following wavenumbers: 1115, 1220, 1254, 1377, 1459, 1565, 1726, 2839, 2868, 2920, and 2952 $\text{cm}^{-1}$.
The NEMS–FTIR spectrum of PVC nanoparticles exhibit vertical lines at the following wavenumbers: 1100, 1254, 1334, 1427, 1564, 2849, 2910, and 2971 $\text{cm}^{-1}$.
B.   NEMS–FTIR spectra of tea bag leachates
The NEMS–FTIR spectra of tea bag leachates (see Figure 5C) exhibit vertical lines at the following wavenumbers: 575, 990, 1116, 1170, 1210, 1235, 1282, 1371, 1437, 1460, 1553, 1642, 1720, 2860, 2930, 3087, and 3300 $\text{cm}^{-1}$.
C.   Derivation of the decadic linear attenuation coefficient for sparse particle film
In the case of sparsely distributed particles with negligible optical interactions, each particle can be considered an independent absorber. For particles much smaller than the wavelength of incident light $(R<<\lambda)$, the absorption cross section is given by [82]:
[…]
where $\tilde{\nu}=1/\lambda$ is the wavenumber, $\tilde{n}(\tilde{\nu})$ the complex refractive index of the particles, and $n_{m}(\tilde{\nu})$ the real refractive index of the surrounding medium.
The particles sit on top of a suspended dielectric film (SiN membrane), which is optically thin compared to the wavelength $\lambda$. Hence, the refractive index of the medium is approximately that of vacuum $n_{m}(\tilde{\nu})\approx 1$, resulting in:
[…]
The decadic volume attenuation coefficient of a particle is given by:
$\mu_{10,v}(\tilde{\nu})=\dfrac{\sigma_{\text{abs}}(\tilde{\nu})}{\ln(10)},$ (10)
which can be converted to the decadic linear attenuation coefficient through division by the particle volume:
[…]
Figure S1: Coffee ring effect. SEM image of (A) 1.25 ng of PP nanoparticles, (B) 1.25 ng of PS nanoparticles, (C) a PS–PP–PVC mixture mixed in a mass ratio of 1:1:1, containing 5 ng of each component. Image also shows a close–up view of the coffee ring structure.
Figure S2: Characterization of PP nanoplastics. NEMS–FTIR spectra of PP nanoplastics deposited on NEMS chips at varying mass loads. The inset highlights the 1377 cm-1 peak, which was used to construct the calibration curve and determine the LoD.
Figure S3: Characterization of PVC nanoplastics. NEMS–FTIR spectra of PVC nanoplastics deposited on NEMS chips at varying mass loads. The inset highlights the 1427 cm-1 peak, which was used to construct the calibration curve and determine the LoD.
Figure S4: Micro- and nanofragments released by tea bags during accelerated aging. (A–C) Magnified views of the NEMS membrane resonator sampled with water in which a tea bag had been soaked for 15 days and exposed to environmental stress.
Figure S5: Correction effects and spectral artifacts in ATR–FTIR. (A) Comparison of non-corrected and corrected ATR–FTIR spectrum of nylon tea bag (TB) material. Corrections applied include "Extended ATR correction" (standard mode) and "Baseline correction" using OPUS software. The figure demonstrates how the applied corrections alter the relative intensities of spectral peaks. (B) Comparison of NEMS-FTIR spectra of 100 nL and 500 nL nylon tea bag leachates (TBL) with ATR-FTIR spectra of 500 nL TBL and the nylon TB material in the 2800–1730 cm-1 region. CO2 peaks and diamond phonon bands are visible in the ATR–FTIR spectra, whereas these spectral artifacts are absent in the NEMS–FTIR spectra.
Figure S6: Experimental setup for the accelerated aging of tea bags. The setup includes a controlled environment chamber for simulating environmental stress through cyclic exposure to elevated temperature and UV radiation, with tea bags immersed in water for the duration of the experiment.
Figure S7: Sampling setups. (A) The PIPEJET nanoDispenser (BioFluidix–Hamilton Freiburg, Germany), equipped with two cameras and an XY stage, enables precise positioning of the NEMS chips beneath the dispensing capillary for accurate droplet deposition onto the membrane. (B) The Drop casting Accessory (Invisible-Light Labs GmbH, Austria) for liquid samples. The accessory allows for precise sample deposition and controlled drying under a mild humidity gradient, ensuring consistent localization of the sample within the perforated membrane area. (C) Schematic representation of the setup, comprising a nebulizer (Portable Aerosol Generation System, PAGS, Handix Scientific Inc., CO, USA) for aerosolizing the liquid into fine droplets, a diffusion dryer (DDU, 570/L, Topas GmbH, Germany), and a NEMS chip holder (Aerosol Flow Adapter, Invisible-Light Labs GmbH, Austria) containing the NEMS chip. The chip holder was placed in an eight-channel filter sampler (FILT, Brechtel Manufacturing Inc., CA, USA), which includes a built–in pump. The sampling flow rate was set to 0.5 L/min for a duration of 5 minutes for each NEMS chip. Sampling is based on the inertial impaction deposition mechanism.
Figure S8: IR spectral features of SiN membranes. Average transmittance and absorptance spectra of 12 individual empty NEMS chips composed of 50 nm thick low–stress SiN. The shaded regions represent the standard deviations associated with each spectrum. The dashed vertical line at 835 $\text{cm}^{-1}$ marks the characteristic SiN peak, where the transmittance and absorptance values, along with their standard deviations, were evaluated.
Figure S9: Dependence of relative photothermal responsivity on the IR beam radius. Data obtained through FEM simulations of a 1x1 mm2 low–stress SiN membrane. The graph highlights how varying the beam radius affects the responsivity of the NEMS membrane due to spatial heat distribution.
Table S1: Deposited and estimated mass of PS NPs. Quantitative comparison of the deposited and estimated mass values for PS nanoparticles on NEMS chips, accompanied by their respective standard deviations. Deposited masses were derived from the dispersion concentrations and droplet volume, with uncertainties propagated from sample preparation (micropipette) and deposition (nanodropper) tools. Estimated masses were calculated using Eq. (6). The main source of uncertainty in estimated mass originates from variability in the measured absorptance at 1452 cm-1.
Deposited mass (ng) SD (ng) Estimated mass (ng) SD (ng)
0.62 0.03 0.8 0.5
1.2 0.1 1.0 0.1
2.5 0.1 2.7 0.2
5.0 0.3 5.5 0.5
10.0 0.5 11 1
20 1 21 1
40 2 43 3